Hello! Can someone please help me solve this problem in a step by step solution?

THanks so much!

Texas-Q Company produces and sells barbeque grills. Texas-Q sells three models: a small portable gas grill, a larger stationa1. What is the sales mix of portable grills to stationary grills to smokers? 2. Compute the break-even quantity of each produ3(6) Prepare an income statement for Texas-Q for the coming year. Refer to the list of Amount Descriptions for the exact word4. Compute the margin of safety for the coming year. The margin of safety for the coming year is $
Answer #1

Answer 1 & 2

Portable Stationary Smokers
Units Sold 15600 52000 5200
Sales Mix 3 10 1
Selling Price 93 203 249
Variable Cost 50 130 144
Contribution Margin 43 73 105
Weighted Average CM                     9.21                      52.14                     7.50
Total Weighted Average CM (A)                  68.86
Total Fixed Cost (B) 2120800
Combined Breakeven Point (B)/(A)                30,800
Breakeven Point Units 6600 22000 2200

Answer 3

Portable Stationary Smokers Total
Units Sold                15,600                    52,000                  5,200
Sales Revenue          14,50,800          1,05,56,000          12,94,800          1,33,01,600
Variable Cost            7,80,000              67,60,000            7,48,800              82,88,800
Contribution Margin            6,70,800              37,96,000            5,46,000              50,12,800
Fixed Cost              21,20,800
Net Operating Income              28,92,000
Contribution margin Ratio = Total Contribution Margin / Total Sales = 5012800 / 13301600 = 37.69%
Breakeven Sales ($) = Fixed Cost / CM (%) = 2120800 / 37.69% = 56,27,600
Portable Stationary Smokers
Bifurcation of Sales 613800 4466000 547800

Answer 4

Total Sales for Coming Year      1,33,01,600
Breakeven Sales          56,27,600
Margin of Safety Sales ($)          76,74,000